OpenStudy (kayders1997):

Can someone please help find the limit of the absolute value of x^2-1 divided by x^2-6x+5 as x approaches 1 from the left

satellite73 (satellite73):

i would factor and cancel first

OpenStudy (kayders1997):

Can you cancel though? Because it is absolute value

satellite73 (satellite73):

don't see why not

OpenStudy (kayders1997):

That's just what my professor told us

satellite73 (satellite73):

you are taking a limit in any case, this thing is undefined at 1, but x cannot be 1

OpenStudy (kayders1997):

Right

OpenStudy (brainzonly):

might help?

OpenStudy (kayders1997):

-1/2???

satellite73 (satellite73):

go ahead and factor and cancel

satellite73 (satellite73):

the limit cannot be negative, that is for sure, as this is never negative

OpenStudy (kayders1997):

Hmmmm

satellite73 (satellite73):

it matters not that they are inside an absolute value, you can still factor and cancel since x is not 1

OpenStudy (brainzonly):

people say you can't factor and cancel absolute value but I've been doing it for years and never got one of those problems wrong. anyway, no one tells you why not

OpenStudy (kayders1997):

So your with absolute value x+1 over (x-5)

satellite73 (satellite73):

yes, now replace x by 1

OpenStudy (kayders1997):

Yeah I got absolute value of 2/-4 so -1/2

satellite73 (satellite73):

absolute value is still there !

OpenStudy (kayders1997):

So 1/2?

satellite73 (satellite73):

yes

OpenStudy (kayders1997):

So it's applied to the whole problem not just the top?

satellite73 (satellite73):

i am not sure what you are asking

OpenStudy (kayders1997):

so it started absolute value of x^2-1 but it didn't have absolute value on the bottom the absolute value applies to the bottom also?

satellite73 (satellite73):

ooh i didn't see that part, lets check

satellite73 (satellite73):

yeah now it is a totally different story

OpenStudy (kayders1997):

That's why I wasn't sure if you could cancel because the bottom doesn't have absolute value and the top does but idk what else you could do algebrically

satellite73 (satellite73):

so lets back up and go slow

OpenStudy (kayders1997):

Ok

satellite73 (satellite73):

first off, is \(x>1\) or is \(x<1\) ?

OpenStudy (kayders1997):

X<1 because it's the left of 1? Right?

satellite73 (satellite73):

yes exactly now we can remove the absolute value sign in the numerator, because if\(x<1\) then \(x^2-1<0\) right ?

OpenStudy (kayders1997):

Yes

satellite73 (satellite73):

in other words \(x^2-1\) is negative that means \[|x^2-1|=1-x^2\] (only since \(x<1\) )

satellite73 (satellite73):

so now the absolute value signs are gone, and you can compute \[\lim{x\to -^1}\frac{1-x^2}{(x-5)(x-5)}\] by cancelling

OpenStudy (kayders1997):

Ok that makes sense!

satellite73 (satellite73):

typo there, i meant \[\lim_{x\to 1^-}\frac{1-x^2}{(x-5)(x-5)}\]

satellite73 (satellite73):

dang another typo!!

OpenStudy (kayders1997):

Lol it's ok!

satellite73 (satellite73):

\[\lim_{x\to 1^-}\frac{1-x^2}{(x-5)(x-1)}\]

satellite73 (satellite73):

now, i believe, the answer will in fact be \(-\frac{1}{2}\) check it

OpenStudy (kayders1997):

Ok

OpenStudy (kayders1997):

Yes that is what I got thank you so much!!!

satellite73 (satellite73):

yw

OpenStudy (triciaal):